1) Consider an economy having a Cobb Douglas production function, where the share of capital income in total income is 1/2. The depreciation rate is delta equals 0.03 , population growth rate is n = 0.02

2) Assume a general savings rate s   , depreciation rate delta    and a production per worker y equals k to the power of alpha   , where 0<alpha  <1. Suppose the savings rate increases. What happens to the golden rule level of capital?

3) Consider an economy that is described by the production function Y equals K to the power of alpha L to the power of 1 minus alpha end exponent   . The depreciation rate is delta    and the savings rate is s. How is the golden rule level of consumption per worker related to the level of consumption per worker obtained using the savings rate s?

QUESTION 2 20 points Save Answer Consider an economy having a Cobb Douglas production tunction, where the share of capital income in total income is 1/2. The depreciation rate is 5 0.03, population growth rate is n 0.02 The golden rule level of captal per worker is The golden rule level of investmeent per worker is The golden rule level of output per worker is The golden rule savings rate is X % where X equals QUESTION 3 10 points Save Answer Assume a general savings rate S ,depreciation rate S and a production per worker y= kwhere 0ca 1. Suppose the savings rate increases. What happens to the golden rule level of capital? Increases decreases remains same indeterminate Save Answer QUESTION 4 10 points Consider an economy that is described by the production tunction Y = KL The depreciation rate is and the savings rate is s. How is the golden rule level of consumption per worken related to the level of consumption per worker obtained using the savings rate s? t depends on the value of s Oless or equal less Oareater enual
Answer #1

Q2) let y = √k, d =.03, n =.02

Then at golden rule level,

MPK = d+n

.5/√k = .05

10 = √k

A)k* = 100

B) i* = sy

= .5*10

= 5

c) y* = 10

d) X = 50

( Golden rule saving rate equals exponent of K in Cobb Douglas)

Q3) option C)

golden rule level capital is independent of saving rate .

Bcoz at golden rule, MPK = d+n

Q4) option D)

At golden rule level, Consumption per capita is maximum

So C*G is greater than or equal to other per capita consumption levels